OpenStudy (brittany16):

Solve the equation. p - 5 - 13p = 150 - 11

OpenStudy (anonymous):

Wolfram Alpha

OpenStudy (anonymous):

p = -12

OpenStudy (anonymous):

http://imgur.com/VDQs9.png

OpenStudy (kamille):

p - 5 - 13p = 150 - 11 p-13p=150-11+5 -12p=144 p=-12